OpenStudy (anonymous):

What is the correct solution to the equation? 5(x – 5) = 15 A. –50 B. –2 C. 8 D. 80

OpenStudy (anonymous):

@jadeybells

OpenStudy (anonymous):

Plug in the numbers given to find your answer (Doesn't teach you to solve for X but it will give you the answer to the question) 5(-50 - 5) = 15 5(-2-5) = 15 5(8-5) = 15 5(80 - 5) = 15

OpenStudy (acxbox22):

no

OpenStudy (anonymous):

Im still not that sure

OpenStudy (acxbox22):